# Investigating Emotions, Social Roles, Well-being and Mobile Usage Behaviors During COVID-19 Home Confinement

### v1

@@ Abstract.
Pandemics and epidemics have been an essential part of human history. It has powerfully shaped human civilization, affecting politics, economics, medicine, and everyday human lives. Home confinement, one of the social isolation restrictions, has proven to be an effective way to prevent a pandemic and significantly reduce the spread of the virus. However, these measures have also led to negative psychological effects, which can be more severe than the physical discomfort caused by the pandemic itself. To address this issue, it is crucial to understand how being confined at home affects people’s psychological states and their use of digital devices. To this end, we conducted an in-situ study with 32 participants living in states affected by COVID-19 lockdowns for three weeks. We examined human emotions, social roles, well-being, and mobile usage behaviors and explored the predictability of affective states using digital traces. Extensive experimental results show that COVID-19 home confinement is associated with decreased valence and increased work time, and that human well-being could be accurately inferred from mobile usage behaviors. We also present a set of interesting findings on how different factors of mobile usage affect human emotions. Overall, the findings of this research have great potential to inform the development of interventions and remote programs to support individuals in similar situations, such as those who are ill or undergoing post-operative home rehabilitation and home work.
Emotion, COVID-19 Lockdown, Home Confinement, Mobile Sensing, Social Roles, Well-being
copyright: acmcopyrightjournalyear: 2023doi: 10.1145/1122445.1122456journal: IMWUTjournalvolume: 37journalnumber: 4article: 52publicationmonth: 8ccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing
@@ 1. Introduction
The coronavirus disease 2019 (COVID-19) pandemic has resulted in a global lockdown, with many people experiencing home confinement for extended periods. Home confinement, defined as the restriction of individuals’ movements and activities within their homes (59), is a common phenomenon that everyone may experience during their lives, whether due to illness or post-recovery (43; 16). Previous research has shown that home confinement may have negative psychological effects, including stress, anxiety, and depression. (6; 73). For instance, Bartoszek et al. (6) found that individuals who were quarantined were associated with high levels of depression, insomnia, loneliness, and everyday fatigue. Understanding the impact of home confinement on mental well-being is essential for developing interventions to support individuals during prolonged periods of confinement.
@@
Existing approaches to measuring human mental well-being during home confinement typically rely on survey-based instruments (73; 28). However, these approaches can be time-consuming and labor-intensive, and may also bring additional burdens to participants, leading to response bias (e.g., reflection bias and social desirability bias (29; 66)). Fortunately, recent advancements in mobile sensing technology offer a promising solution for predicting human mental well-being (68; 20). For instance, Rashid et al. (68) modeled social anxiety by utilizing sparsely collected mobile sensor data from socially anxious participants, while Chowed et al. (20) measured depression, social anxiety, affect, and social isolation of undergraduate students using GPS signals collected through mobile devices. However, most sensing technologies are designed to monitor daily lives in general, rather than specifically for home confinement situations. As such, the patterns and daily behaviors captured by these technologies may differ significantly from those experienced during home confinement.
@@
Studying the in-situ mental well-being and mobile usage behaviors of individuals confined to their homes can be challenging due to the variability of their degree of confinement, which may affect the accuracy and validity of collected data. Nevertheless, the COVID-19 lockdown presents a unique opportunity to study human well-being during home confinement. On one hand, the strict measures put in place have created similar confinement conditions for all participants due to policy mandates. On the other hand, the COVID-19 home confinement has led to an increasing reliance of digital technologies(64), with srategies such as stay-at-home and work-from-home being recommended worldwide (69). As a result, people are increasingly turning to smartphones for social networks, online meetings, Internet surfing, or video games (69). Recent research shows that COVID-19 lockdowns have caused around 70% of Internet users to increase their smartphone usage (77). The reliance on technology during the COVID-19 lockdown has provided an excellent way to understand people’s mental well-being during home confinement situations.
@@
In this research, we aim to explore the relationship between people’s mental well-being and their interactions with mobile devices during COVID-19 home confinement. Specifically, we are focusing on emotions and social roles because these aspects have a significant impact on overall well-being (22; 61). To assess emotional states, we utilize the widely accepted valence and arousal dimensions within the Circumplex Model of Affect (74) to assess emotional states. Social roles, as defined by Role Theory (8), are associated with norms, expectations, rules and behaviors that provide valuable information about the user’s current context, enabling us to understand their status, task boundaries and work-life balance. This, in turn, can have a significant impact on an individual’s overall well-being and mental health. Furthermore, analyzing app usage records and extracting user activities from them allows for a more granular understanding of user behavior. As a result, our research questions are:
@@
To answer the research questions, we conducted a three-week data collection during the COVID-19 lockdown in Melbourne, Australia. To answer RQ1, we analyzed patterns of perceived valence, arousal and social roles during COVID-19 home confinement using Experience Sampling Method (ESM) (7). We also investigated the influence of the COVID-19 lockdown on emotions, social roles, and work-life balance by examining the end-of-day survey data. To address RQ2, we proposed the App-based Information Gain Temporal Segmentation (AIGTS) to extract user activities (e.g., video-watching, social, deep work) from app usage records by identifying segments and transition times between apps during the day. Finally, we demonstrated the predictability of perceived well-being from smartphone usage behaviors for RQ3. All scripts for this study are open source. 111link to be shared after the double-blind review. In summary, our contributions are as below:
@@
  • We collected a comprehensive dataset from 32 participants over a period of 3 weeks during COVID-19 home confinement in Melbourne, Australia. The participants completed ESM questionnaires and end-of-day surveys to report their emotions, social roles, well-being, etc. Their smartphone and desktop usage were recorded through an app running in the background. The resulting dataset includes 502,485 records of users’ context and phone usage, along with 1,749 ESM responses and 265 end-of-day survey responses.
@@
  • We conducted an in-depth analysis of the participants’ perceived valence, arousal, and social roles during the COVID-19 home confinement, and explore how lockdown affects people’s well-being and work-life balance. We discovered that the lockdown tended to reduce valence and increase work hours due to interruptions or blurring of personal and work boundaries.
@@
  • We proposed an unsupervised activity recognition method called App-based Information Gain Temporal Segmentation (AIGTS). This method enables the extraction of user activities from app usage records. By identifying different segments of app sequences and applying clustering techniques, we were able to discover 16 distinct user activities in an unsupervised manner. We also discussed the correlations between these user activities and human well-being.
@@
  • We conducted extensive experiments to predict human well-being based on smartphone usage behaviors. Experimental results demonstrated the superior prediction performance of the proposed model on mental well-being. Additionally, we derived different factors that contribute to understanding well-being and discussed the top predictors of human emotion during home confinement.
The remainder of the paper is as below. Section  2 presents related work on the impact of COVID-19 home confinement on human well-being, as well as recent advances in mobile sensing technologies for predicting well-being. Section  3 describes the data collection procedures. In Section 4, we explore patterns of human well-being during COVID-19 home confinement and investigate how the lockdown affects emotion and work-life balance. Next, we extract user activities through app sequences and explore correlations between user activities and human well-being in Section 5. Then, in Section  6, we present experimental results for predicting mental well-being during home confinement and provide a detailed discussion of the findings. Section 7 discusses the implications and Section 8 lists the limitations of our work. Finally, we summarize this research in Section  9 and indicate the potential direction of future work.
@@ 2.1. Impact of the Pandemic Restrictions on Human Well-being
Pandemics and epidemics have played an essential role throughout human history (82). In the last century alone, there were Spanish flu (1918-1920), Asian flu (1957-1958), AIDS (1981-present day), H1N1 Swine flu (2009-2010), Ebola (2014-2016), COVID-19 (2019-present day), affecting people all over the world. Researchers explored the psychological symptoms arising from pandemics and epidemics, and found that restrictive measures (e.g., social distancing, quarantine and isolation) were associated with negative psychological effects (82; 19; 11), including post-traumatic stress symptoms (71), depression (32), stress (25), emotional disturbance (92), irritability(44), low mood (44), confusion (15), anger (52) and anxiety-induced insomnia (25). Even after the quarantine ends, the psychological effects of quarantine can still be observed months or years later (38).
According to the World Health Organization  (89), the COVID-19 pandemic has severally impacted human well-being. The potential for physically sick or discomfort from COVID-19 is considered less of concern than the psychological and social impact (35). A growing body of literature has explored the impact of the COVID-19 on emotional and mental health (90; 21; 45; 42). Daly et al.  (21) explored the effects of COVID-19 on mental health in the UK. They found a significant increase in observed mental health problems, particularly for those aged 18-34 and women. Fountoulakis et al. (26) found that, during the COVID-19 lockdown in Greece, anxiety increased in more than 45% and suicidal thoughts increased in 10.40%. Khubchandani et al. (41) found the prevalence of depression (39%), anxiety (42%) and psychological distress (39%) after initial COVID-19 lockdown in the USA. The rate of serious mental health issues in the USA has more than doubled compared to pre-COVID-19 rates in the year 2019. Killgore et al. (42) found that loneliness has increased during the COVID-19 pandemic, especially for those under lockdown restrictions.
Despite the existing research on the impact of pandemic restrictions on human well-being, much of it is limited to specific regions and conducted during different phases of varying degrees of restrictive measures. This makes it challenging to fully comprehend the effects of strict home confinement on individuals’ well-being. In light of this, the first research question (RQ1) focuses on studying the emotions, social roles and work-life balance change during COVID-19 home confinement, specifically in metropolitan Melbourne where the strictest restrictive measures (Stage 4) were implemented.
@@ 2.2. Impact of the Pandemic Restrictions on Smartphone Usage
The outbreak of COVID-19 has interrupted normal activities and caused many organisations to close, forcing people to work from home employing technology (84). Twitter even made an announcement to make working remotely a permanent option (12). As people perform their duties by staying at home or working from home, they increasingly turned to digital devices to communicate, interact and conduct job responsibilities (69). Compared to pre-lockdown, Internet service usage has increased from 40% to 100%, and usage of video conferencing apps such as Zoom increased tenfold (77; 64). A recent global study (77) reported that around 70% of internet users used their smartphones or mobile phones more often as a direct result of the COVID-19 lockdown, suggesting that the COVID-19 lockdown policies may lead to the overuse of smartphones (69).
Smartphone overuse may affect the physical and psychological health of users, such as neck/lower back/shoulder pain, anxiety and depression (69; 13; 4). Ratan et al. (69) pointed out that problematic smartphone usage affects the ‘reward system’ of the brain and subsequent harms will persist in the form of various emotional disorders even after the lockdown effects are lifted. Arora et al. (4) explored the problematic phone use and its impact on mental health during the COVID-19 lockdown, and found that ‘Sadness’ and ‘anger’ were the dominant emotions in addictive Twitter users, while ‘valence’ was the dominating emotion among non-addictive users. Busch et al. (13) revealed that excessive mobile phone use may severely impact people’s mental health and human well-being. Especially during the COVID-19 lockdown, people’s mobile phone usage differs than normal, which increases the risk of overusing smartphones and affects their mental health (5).
@@ 2.3. Investigating User Activities from App Sequences
By analysing app sequences, researchers have been able to gain insights into various aspects of individuals’ behaviours in daily lives, including social roles  (2) and interruptiblity (62; 63). Deploying the approach ‘Application as a Sensor’, Anderson et al.  (2) found that app sequences on smartphones are correlated with individuals’ private and work social roles. Similarly, Okoshi et al. (62; 63) explored app sequences data and applied machine learning techniques to predict when a break-point in activity occurs for intelligent notification delivery. These approaches focus on identifying breakpoints between apps as a means to understand interruptibility. However, none of the previous research studies have utilised breakpoints to identify activity segments and predict contextual variables (i.e., user activity) from these segments.
Moreover, previous studies often relied on simplistic methods of aggregating app usage, such as calculating the total amount of time an app is used over a specific period. However, those approaches fail to capture the nuances and transitions between different app activities, limiting understanding of user behaviours. Moreover, traditional self-report surveys may impose a burden on users, potentially resulting in biased or incomplete data. Conversely, analyzing app sequences in an unsupervised manner can yield a wealth of information pertaining to user behaviors and mental well-being. With this in mind, our research question RQ2 aims to explore the unsupervised extraction of user activity from app usage records.
@@ 2.4. Inferring Human Well-being Using Mobile Sensing Technologies
The advent of sensing technology has opened up new avenues for understanding and predicting human behaviours (51) and psychological stages, e.g., valence (37), mood (56; 93; 55), engagement (30; 36), social anxiety (20; 68), depression (86; 83), etc. Most studies integrate data from various sources for building comprehensive prediction models. For instance, Gao et al. (30) predicted emotional, cognitive and behavioral student learning engagement in an Australian high school using wearable sensors and environmental sensors. Wang et al. (86) measured depression of undergraduate students at Dartmouth College using passive mobile sensing data and wearable data. Jaques et al. (37) modeled the valence of MIT undergraduate students using physiological signals collected from wristbands and behavior data collected from mobile devices. Morshed et al. (56) presented to predict mood instabilities using passively sensed smartphone and wearable signals.
@@
While utilizing multiple sensing sources can improve the accuracy of psychological state prediction, there is a growing interest in using mobile sensing data alone due to its convenience and ability to provide rich information about an individual’s behavior, e.g., physical activity, communication patterns, and location. For example, Hintze et al. (34) analyzed mobile device usage characteristics such as session length, interaction frequency and daily usage in locked/unlocked state with respect to location context and diurnal pattern, to gain insights into how people use their smartphones and tablets. Madan (51) et al. designed a mobile sensing platform and found that location and communication logs could be used to model individual symptoms, long-term health outcomes, and diffusion of opinions in society.
@@
Some studies demonstrated the potential of mobile sensing data alone for understanding psychological states, e.g., anxiety (68; 83; 20), mood (93; 55), emotion (81), social roles (2) and interruptibility (62; 63). Zhang et al. (93) proposed an Android app to detect compound emotion utilizing smartphone usage data. Meegahapola et al. (55) explored the effect of geographical diversity on mood inference. They examined eight countries and compared the continent-specific, country-agnostic, and multi-country approaches trained for two mood inference tasks. Tlachac et al. (83) collected smartphone call logs and text logs during the COVID-19 pandemic for developing machine learning models to predict anxiety and depression. Tag et al. (81) developed an Andriod app to measure emotion trajectories and examine how digital emotion regulation unfolds in naturalistic settings. Rashid et al. (68) utilized sparsely collected mobile sensor data to predict perceived social anxiety. Chowed et al. (20) explored the fine-grained GPS data collected through mobile devices to model depression, social anxiety, affect, and social isolation.
@@
While numerous studies have utilized mobile sensing to predict human well-being, there is a significant research gap in the specific context of home confinement. Unlike daily life situations, individuals’ behaviors may differ considerably during home confinement, such as increased reliance on digital devices, altered perception of time, and remote working. Although a few studies have examined psychological states using mobile sensing during the COVID-19 pandemic (83; 60), these works largely focus on broader environments rather than strictly home confinement. Investigating the unique characteristics of home confinement could provide valuable insights for understanding individuals in similar situations such as post-operative recovery, parental leave, aging or disability-related limited mobility. Therefore, we aim to address RQ3 by exploring the possibility of modeling emotions during home confinement and identifying significant features for prediction.

### v2

# ”Living Within Four Walls”: Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement

@@ Abstract.
Home confinement, a situation experienced by individuals for reasons ranging from medical quarantines, rehabilitation needs, disability accommodations, and remote working, is a common yet impactful aspect of modern life. While essential in various scenarios, confinement within the home environment can profoundly influence psychological well-being and digital device usage. In this study, we delve into these effects, utilising the COVID-19 lockdown as a special case study to draw insights extending to various homebound situations. We conducted an in-situ study with 32 participants living in states affected by COVID-19 lockdowns for three weeks and analysed their emotions, well-being, social roles, and mobile usage behaviours. We extracted user activity from app usage records in an unsupervised manner, and experimental results revealed that app usage behaviours are effective indicators of emotional well-being in confined environments. Our research has great potential for developing supportive strategies and remote programs, not only for people facing similar medical isolation situations, but also for individuals in long-term home confinement, such as those with chronic illnesses, recovering from surgery, or adapting to permanent remote work arrangements.
Emotion, Home Confinement, Mobile Sensing, COVID-19 Lockdown, Social Roles, Well-being
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.XXXXjournal: TOCHIccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing
@@ 1. Introduction
Home confinement, defined as the restriction of individuals’ movements and activities within their homes (63), is a common phenomenon that everyone may experience during their lives, whether due to illness or post-recovery (47; 17). Previous research indicated that home confinement may adversely affect mental well-being, leading to stress, anxiety, and depression (7; 78). For instance, Bartoszek et al. (7) found that individuals who were quarantined were associated with high levels of depression, insomnia, loneliness, and everyday fatigue. Understanding the impact of home confinement on mental well-being is essential for developing effective interventions to support individuals during prolonged periods of confinement.
@@
Existing approaches to measuring human mental well-being during home confinement typically rely on survey-based instruments (78; 30). However, these approaches can be time-consuming and labour-intensive, and may also bring additional burdens to participants, leading to response bias (e.g., reflection bias and social desirability bias (32; 71)). Fortunately, recent advancements in mobile sensing technology offer a promising solution for predicting human mental well-being unobtrusively (73; 21). For instance, Rashid et al. (73) modelled social anxiety by utilizing sparsely collected mobile sensor data from socially anxious participants, while Chowed et al. (21) measured depression, social anxiety, affect, and social isolation of undergraduate students using GPS signals collected through mobile devices.
In home confinement scenarios, traditional mobile sensing methods, heavily relying on sensors like accelerometers and GPS to track physical behaviours, encounter unique challenges. These methods, which infer mental well-being from patterns such as visited locations, steps taken, and stationary periods (73), become less effective when individuals are confined at home. During home confinement, individuals’ interaction with their smartphones changes significantly. Rather than carrying their devices with them, people tend to leave their phones in one place. This shift in usage patterns diminishes the effectiveness of conventional metrics such as step counts and travel patterns, which are less indicative of mental well-being in a stationary context. Consequently, there is a shift towards analyzing more nuanced data types, such as application usage patterns, to better understand well-being under home confinement. Therefore, in this research, we focus on these finer-grained user activities to uncover deeper insights into the subtle changes in mental well-being during home confinement.
@@
Another challenge in studying the in-situ mental well-being of individuals confined to their homes is establishing a natural yet controlled environment. Home confinement situations vary significantly among individuals, influenced by different reasons and individual differences, making effective modelling challenging. Nevertheless, the COVID-19 lockdown presents a unique opportunity to study human well-being during home confinement. The strict measures in place have created similar confinement conditions for all participants due to policy mandates. Additionally, the COVID-19 home confinement has led to an increasing reliance on digital technologies (68), with strategies such as stay-at-home and work-from-home being recommended worldwide (74). Recent research indicates that COVID-19 lockdowns have caused around 70% of Internet users to increase their smartphone usage (82). This reliance on technology provides an excellent way to understand people’s mental well-being during home confinement.
@@
In this research, we aim to explore the relationship between people’s mental well-being and their interactions with mobile devices during home confinement. Specifically, we focus on emotions and social roles, as these aspects significantly impact overall well-being (23; 65). To assess emotional states, we utilize the widely accepted Valence and Arousal dimensions within the Circumplex Model of Affect (79) to assess emotional states. Social roles, as defined by Role Theory (9), involve norms, expectations, rules and behaviours that provide valuable information about the user’s current context, enabling us to understand their status, task boundaries and work-life balance. This, in turn, can have a significant impact on an individual’s overall well-being and mental health. Furthermore, analyzing app usage records and extracting user activities from them allows for a more granular understanding of user behaviours. As a result, our research questions are:
@@
To answer the research questions, we conducted a three-week data collection during the COVID-19 lockdown in Melbourne, Australia. To answer RQ1, we analyzed patterns of perceived valence, arousal and social roles during COVID-19 home confinement using Experience Sampling Method (ESM) (8). We also investigated the influence of the COVID-19 lockdown on emotions, social roles, and work-life balance by examining the end-of-day survey data. To address RQ2, we proposed the App-based Information Gain Temporal Segmentation (AIGTS) to extract user activities (e.g., video-watching, social, deep work) from app usage records by identifying segments and transition times between apps during the day. Finally, we demonstrated the predictability of perceived well-being from smartphone usage behaviours for RQ3. All scripts for this study are open source. 111link to be shared after the double-blind review. In summary, our contributions are as below:
@@
  • We collected a comprehensive dataset from 32 participants during a 3-week COVID-19 home confinement period in Melbourne, Australia. The participants completed ESM questionnaires and end-of-day surveys to report their emotions, social roles, well-being, etc. Their smartphone and desktop usage were recorded through an app running in the background. The resulting dataset includes 502,485 records of users’ context and phone usage, along with 1,749 ESM responses and 265 end-of-day survey responses.
@@
  • We analyzed participants’ perceived valence, arousal, and social roles during home confinement, and explored how lockdown affected well-being and work-life balance. We discovered that the lockdown tended to reduce valence and increase work hours due to interruptions or blurring of personal and work boundaries.
@@
  • We proposed an unsupervised activity recognition method, App-based Information Gain Temporal Segmentation (AIGTS), to extract user activities from app usage records. This approach autonomously identified 16 distinct user activities in an unsupervised manner, and we then explored the correlations between these user activities and human well-being.
@@
  • We conducted extensive experiments to predict human well-being in confined environments from smartphone usage behaviours, demonstrating the superior performance of our model. Additionally, we identified key factors contributing to well-being and discussed the primary predictors of human emotion during home confinement.
@@ 2.1. Impact of the Pandemic Restrictions on Human Well-being
Pandemics and epidemics have played an essential role throughout human history (87). In the last century alone, there were Spanish flu (1918-1920), Asian flu (1957-1958), AIDS (1981-present day), H1N1 Swine flu (2009-2010), Ebola (2014-2016), COVID-19 (2019-present day), affecting people all over the world. Researchers explored the psychological symptoms arising from pandemics and epidemics, and found that restrictive measures (e.g., social distancing, quarantine and isolation) were associated with negative psychological effects (87; 20; 12), including post-traumatic stress symptoms (76), depression (35), stress (26), emotional disturbance (97), irritability(48), low mood (48), confusion (16), anger (56) and anxiety-induced insomnia (26). Even after the quarantine ends, the psychological effects of quarantine can still be observed months or years later (42).
According to the World Health Organization  (94), the COVID-19 pandemic has severely impacted human well-being. The potential for physically sick or discomfort from COVID-19 is considered less of a concern than the psychological and social impact (38). A growing body of literature has explored the impact of COVID-19 on emotional and mental health (95; 22; 49; 46). Daly et al.  (22) explored the effects of COVID-19 on mental health in the UK and found a significant increase in observed mental health problems. Fountoulakis et al. (28) found that, during the COVID-19 lockdown in Greece, anxiety increased in more than 45% and suicidal thoughts increased by 10.40%. Khubchandani et al. (45) found the prevalence of depression (39%), anxiety (42%) and psychological distress (39%) after the initial COVID-19 lockdown in the USA. The rate of serious mental health issues in the USA has more than doubled compared to pre-COVID-19 rates in the year 2019. Killgore et al. (46) found that loneliness has increased during the COVID-19 pandemic, especially for those under lockdown restrictions.
Despite the existing research on the impact of pandemic restrictions on human well-being, much of it is limited to specific regions and conducted during different phases of varying degrees of restrictive measures. This makes it challenging to fully comprehend the effects of strict home confinement on individuals’ well-being. In light of this, RQ1 focuses on studying the emotions, social roles and work-life balance change during COVID-19 home confinement, specifically in metropolitan Melbourne where the strictest restrictive measures (Stage IV) were implemented.
@@ 2.2. Impact of the Pandemic Restrictions on Smartphone Usage
The outbreak of COVID-19 has interrupted normal activities and caused many organisations to close, forcing people to work from home employing technology (89; 27; 4). Twitter even made an announcement to make working remotely a permanent option (13). As people perform their duties by staying at home or working from home, they increasingly turn to digital devices to communicate, interact and conduct job responsibilities (74). Compared to pre-lockdown, Internet service usage has increased from 40% to 100%, and usage of video conferencing apps such as Zoom increased tenfold (82; 68). A recent global study (82) reported that around 70% of internet users used their smartphones or mobile phones more often as a direct result of the COVID-19 lockdown, suggesting that the COVID-19 lockdown policies may lead to the overuse of smartphones (74).
Smartphone overuse may affect the physical and psychological health of users, such as neck/lower back/shoulder pain, anxiety and depression (74; 14; 5). Ratan et al. (74) pointed out that problematic smartphone usage affects the ‘reward system’ of the brain and subsequent harms will persist in the form of various emotional disorders even after the lockdown effects are lifted. Arora et al. (5) explored the problematic phone use and its impact on mental health during the COVID-19 lockdown, and found that ‘sadness’ and ‘anger’ were the dominant emotions in addictive Twitter users, while ‘valence’ was the dominating emotion among non-addictive users. Busch et al. (14) revealed that excessive mobile phone use may severely impact people’s mental health and human well-being. Especially during the COVID-19 lockdown, people’s mobile phone usage differs than normal, which increases the risk of overusing smartphones and affects their mental health (6).
@@ 2.3. Investigating User Activities from App Sequences
By analysing app sequences, researchers have gained insights into various aspects of individuals’ daily behaviours, including social roles  (2) and interruptiblity (66; 67). Deploying the approach ‘Application as a Sensor’, Anderson et al.  (2) found that app sequences on smartphones are correlated with individuals’ private and work social roles. Similarly, Okoshi et al.  (66; 67) utilized app sequence data and machine learning techniques to predict breakpoints in activity for intelligent notification delivery. These studies focus on identifying breakpoints between apps to understand interruptibility. However, no previous research has used breakpoints to identify activity segments and predict contextual variables (i.e., user activity) from these segments.
Moreover, previous studies often relied on simplistic methods of aggregating app usage, such as calculating the total amount of time an app is used over a specific period. However, those approaches fail to capture the nuances and transitions between different app activities, limiting understanding of user behaviours. Moreover, traditional self-report surveys may impose a burden on users, potentially resulting in biased or incomplete data. Conversely, analyzing app sequences in an unsupervised manner can yield a wealth of information about user behaviours and mental well-being. With this in mind, RQ2 aims to explore the unsupervised extraction of user activity from app usage records.
@@ 2.4. Inferring Human Well-being Using Mobile Sensing Technologies
The advent of sensing technology has opened up new avenues for understanding and predicting human behaviours (55) and psychological stages, e.g., valence (41), mood (60; 98; 59), engagement (33; 39; 31), social anxiety (21; 73), depression (91; 88), etc. Most studies integrate data from various sources for building comprehensive prediction models. For instance, Gao et al. (33) predicted emotional, cognitive and behavioural student learning engagement in an Australian high school using wearable sensors and environmental sensors. Wang et al. (91) measured depression of undergraduate students at Dartmouth College using passive mobile sensing data and wearable data. Jaques et al. (41) modeled the valence of MIT undergraduate students using physiological signals collected from wristbands and behaviour data collected from mobile devices. Morshed et al. (60) presented to predict mood instabilities using passively sensed smartphone and wearable signals.
@@
While utilizing multiple sensing sources can improve the accuracy of psychological state prediction, there is a growing interest in using mobile sensing data alone due to its convenience and ability to provide rich information about an individual’s behaviour, e.g., physical activity, communication patterns, and location. For example, Hintze et al. (37) analyzed mobile device usage characteristics such as session length, interaction frequency and daily usage in locked/unlocked state with respect to location context and diurnal pattern, to gain insights into how people use their smartphones and tablets. Madan (55) et al. designed a mobile sensing platform and found that location and communication logs could be used to model individual symptoms, long-term health outcomes, and diffusion of opinions in society.
@@
Some studies demonstrated the potential of mobile sensing data alone for understanding psychological states, e.g., anxiety (73; 88; 21), mood (98; 59), emotion (86), social roles (2) and interruptibility (66; 67). Zhang et al. (98) proposed an Android app to detect compound emotion utilizing smartphone usage data and Meegahapola et al. (59) explored the effect of geographical diversity on mood inference. Tlachac et al. (88) collected smartphone call logs and text logs during the COVID-19 pandemic for developing machine-learning models to predict anxiety and depression. Tag et al. (86) developed an Andriod app to measure emotion trajectories and examine how digital emotion regulation unfolds in naturalistic settings.
@@
While numerous studies have utilized mobile sensing to predict human well-being, there is a significant research gap in the specific context of home confinement. Unlike daily life situations, individuals’ behaviours may differ considerably during home confinement, such as increased reliance on digital devices, altered perception of time, and remote working. Although a few studies (88; 64) have examined psychological states using mobile sensing during the COVID-19 pandemic, these works largely focus on broader environments rather than strictly home confinement. Investigating the unique characteristics of home confinement could provide valuable insights for understanding individuals in similar situations such as post-operative recovery, parental leave, ageing or disability-related limited mobility. Therefore, we aim to address RQ3 by exploring the possibility of modelling emotions during home confinement and identifying significant features for prediction.